Operations Report – December 22nd

Crew 236 Operations Report 22Dec2021

SOL: 3

Name of person filing report: Pavithra Ravi

Non-nominal systems: Suits and Radios — a headpiece is malfunctioning

Notes on non-nominal systems: One of the headsets is malfunctioning. It keeps transmitting even when the earpiece button is not pressed, resulting in loud interference. It was verified that the issue is with the headset and not the radio (the issue persisted when the radio was swapped out).

Generator:

Hours run: 13 hours 18 minutes

From what time last night: 18:27

To what time this morning: 07:45

List any additional daytime hours when the generator was run: None

Solar— SOC % (Before generator is run at night): Information provided by Outpost

Station Propane Reading – % Information provided by Outpost

Ethanol Free Gasoline: Not in use

Water (loft tank): 54 gallons

Water Meter: 1539651 units

Water (static tank): approx. 370 gallons

Static to Loft Pump used: Yes. The loft tank was refilled from 12.5 to 54 gallons.

Water in GreenHab: approx. 222 gallons

Water in ScienceDome: N/A

Toilet tank emptied: No

Deimos rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours (before EVA 1): N/A

Beginning charge (Before EVA 1): N/A

Ending charge (On return from EVA 1, before recharging): N/A

Hours (before EVA 2): N/A

Beginning charge (Before EVA 2): N/A

Ending charge (On return from EVA 2, before recharging): N/A

Currently charging: Yes

Opportunity rover used: Yes

Hours: 76.4

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 154.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: None

ATV’s Used (Honda, 300, 350.1, 350.2, 350.3): None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: None

Summary of internet: Nominal

Summary of suits and radios: One of the headsets is malfunctioning. It keeps transmitting even when the earpiece button is not pressed, resulting in loud interference. It was verified that the issue is with the headset and not the radio (the issue persisted when the radio was swapped out).

Summary of Hab operations: Nominal

Summary of GreenHab operations: Nominal

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: In correspondence with the
astronomy support team.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support:

The heater filter was changed today. Should the old filter be added to the non-burnables, or be placed in the back airlock for removal?

Any suggestions on how to fix the earpiece? If it is not fixable,
could a new earpiece be provided so each crew member can have their own?

Sol Summary – December 22nd

Crew 236 Sol Summary Report 22-DEC-2021

Sol: 3

Summary Title: Naps on Mars?

Author’s name: Kasey Hilton

Mission Status: Nominal

Sol Activity Summary: Life on Mars? *Cue David Bowie* We don’t know. Naps on Mars? Absolutely. The crew had an early wake-up and a quick breakfast before heading out to on EVA.

Due to a malfunction in comms, the EVA was cut short. The afternoon was spent relaxing and recharging, with 5 out of the 7 crew members taking naps. A bit of work was squeezed in before ending the evening with jambalaya and card games.

Look Ahead Plan: Tomorrow, Cesare (XO), Ben (JOU), Tyler (SCI), and Pavi (ENG) will go on a second EVA. The remaining crew members will stay at the Hab and continue to work on research and outreach.

Weather: Partly overcast, temperature in the 30s.

Crew Physical Status: A resounding “WOOHOO!” followed by a few thumbs up from the crew. Everyone is still in good health and spirits.

EVA: Today’s EVA ended earlier than planned due to a radio malfunction that caused a crew member to lose comms. Tomorrow the same EVA will be attempted again in order to complete the desired research.

Reports to be filed: Operations Report, GreenHab Report, EVA Request, EVA Report, Health and Safety Report, Sol Summary, and Journalist Report (with photos), Astronomy Report

Support Requested: Any information about the James Webb launch would be great (if it’s been delayed further, when is the launch scheduled for, etc.)

Operations Report Dec 07th – 12th

MDRS Operations Report 7-12- Dec-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Electrician will be here either Monday or Tuesday

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours:

Average 50.08

Minimum 47.10

Maximum 58.40

Notes on solar: We are adjusting power use to compensate for the weather. We had two days of overcast weather and SOC got to 16% but on sunny days and with our conservation efforts, we are still running on solar alone.

Propane Reading, station tank – 59 % %

Propane Reading, director tank— 66 %

Propane Reading, intern tank— 38%

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 400

Static to Loft Pump used – yes

Water in GreenHab – 250 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 500 gallons

Hab toilet tank emptied: yes

Perseverance rover used: yes

Hours: 225.4

Beginning charge: 100

Ending charge: 69

Currently charging: no

Sojourner rover used: yes

Hours: 158.1

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: Yes

Hours: 145.0

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: No

Hours: 75.6

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: Yes

Hours: 154.9

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Percy was rolled onto its driver’s side on Saturday at Robert’s Rock Garden. A tow company was called to upright it and a damaged rim was fixed enough for us to get the rover back to the station. Current assessed damage is the front driver’s tire rim, which will need to be replaced, and structural damage to the plastic outer parts. There may be some readjustment required for the bed. The batteries and “engine” appear fine, the brakes and gears are operational and it drives normally. Further assessment will be done in the next few days and any other damage will be reported in our next report. Percy has been removed from the station and is at the Outpost pending repairs.

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town to get water

CrewCar used and why, where? No. Battery will be replaced early this wek.

Luna used and why, where? Yes, around MDRS

General notes and comments: Nothing to report

Summary of internet: Crew internet was down for a day but has been restored. All other accounts nominal.

Summary of suits and radios: All unplugged except for communication radios

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Hab is fully operational with 4 people staying there. Both the furnace and the wall heater are being used for heat. This way we can conserve power overnight as the wall heater does not use power.

Summary of Outpost operations: Work continues on the Intern Trailer for a move in date of December 18th. Currently the water system is frozen as the wall heater was not on overnight. We will inspect for broken pipes as soon as the water thaws.

Summary of GreenHab operations: The new LED lights have been installed and we will be using supplemental lighting again beginning tonight from 7 pm to midnight.

Summary of ScienceDome operations: Used for science teaching during the recent three day visit with Ogden Prep.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Observatory was used for a short time during the class visit. There is currently a mount problem with the Robotic Observatory.

Summary of health and safety issues: Minor injuries reported during rover rollover.

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report Dec 04th – 08th

MDRS Operations Report 4-8- Dec-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours:

Average 49.73

Minimum 44.60

Maximum 58.40

Notes on solar: We are adjusting power use to compensate for the weather.

Propane Reading, station tank – FULL %

Propane Reading, director tank— FULL %

Propane Reading, intern tank— 50 %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 500 gallons

Static to Loft Pump used – yes

Water in GreenHab – 250 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 550 gallons

Hab toilet tank emptied: no

Perseverance rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Sojourner rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Spirit rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Opportunity rover used: No

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: No

Curiosity rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Notes on rovers: Percy should be used to short range EVA’s only as the batteries are not at strong as the remaining rovers, which all have new batteries.

ATV’s Used: (Honda, 350.1, 350.2, 300): The tires have been successfully mounted on the rear of the Honda and it is once again operational.

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town to get water

CrewCar used and why, where? Yes, To Grand Junction. Battery does not have enough cranking power in the below freezing mornings.

Luna used and why, where? Yes, to Loa for supplies for the astrobiology project

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal

Summary of suits and radios: All unplugged except for communication radios

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Hab is fully operational with 4 people staying there.

Summary of Outpost operations: We have been working on the intern trailer. Painting was completed today, except for the area around the bunks and the bunks themselves. That will need to be done once there are fewer people so the bunk paint can harden before use. New mattress installed on the lower bunk, curtain installed between bathroom and bedroom. Carpet removed in the living area. Tomorrow we will complete the cleanup and the interns and Atila will move in on Thursday morning.

Summary of GreenHab operations: Danielle has taken charge of the GreenHab.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report Nov 27th – Dec 03rd

MDRS Operations Report 30-NOV-2021- 3 Dec-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Charge controller problem resolved

Action taken for non-nominal systems: I was able to reset on the second try

Generator: The electrician who wired the system together originally will be coming to figure out the problem on either the 13th or 14th. We had a rather intimidating experience with the system and I didn’t want us to get hurt with my experiments. I can’t fix this one.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours:

Average 50.67

Minimum 45.90

Maximum 58.40

Notes on solar: These numbers are reasonable with the addition of the heat in the Hab and lights in the GreenHab,

Propane Reading, station tank – FULL %

Propane Reading, director tank— FULL %

Propane Reading, intern tank— 50 %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 300

Static to Loft Pump used – yes

Water in GreenHab – 100 gallons (not accessible)

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 150 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: No

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: No

Opportunity rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: No

Curiosity rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: We are unplugging the rovers at night on a rotating basis and only charging them in the daytime

ATV’s Used: (Honda, 350.1, 350.2, 300): Honda’s new tires came but one of them went flat and was returned and fixed. We should get them on and the machine operational tomorrow.

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town to get water and run errands.

CrewCar used and why, where? Yes, to Green River to return Allan to the station and to Bicknell for me to get my COVID booster

Luna used and why, where? Yes, to pick up Atila, Allan and Sebastian in Green River

General notes and comments: CrewCar needs a new battery before the crews return but is fine when run everyday. The tow lights have disconnected from the CrewCar hitch and will be repaired tomorrow.

Summary of internet: All accounts are nominal

Summary of suits and radios: All unplugged except for communication radios

Campus wide inspection, if action taken, what and why? : Nothing to report

Summary of general operations: All propane tanks have been switched and we are awaiting pick up of the old tanks, which are currently stored next to the Hab tank.

Summary of Hab operations: Atila and Sebastian are crewing the Hab while we finish the intern trailer. We have two more interns arriving this weekend and they will stay there too.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Plants were fertilized. Supplemental light with old fixtures from 7 pm-midnight.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.